# Examining Racial Stereotypes in YouTube Autocomplete Suggestions

### v1

@@ 4.3 Culture
First, we found autocompletes that indicated appropriation of cultural practices, highlighting asymmetric power relations between racial groups. For example, several autocompletes related to White groups coopted indigenous Black cultures while belittling its native contexts (e.g., “White girl can dance African”). Other autocompletes revealed how White people could appropriate specific aspects of popular Asian culture, e.g., “Can a White guy win an Asian beauty pageant?,” “Can a White guy make a K-pop group?.” The keyword ‘Black’ also elicited one autocomplete suggestion related to the adoption of Asian cultural elements (“Can a Black person become a K-pop idol?”).
Second, autocompletes associated with culturally specific attire and beauty standards also indicated instances of cultural appropriation, where members of the majority group coopted traditional garments, hairstyles, or makeup practices of other minority groups. For example, autocompletes related to garments or hairstyles for White groups included traditional styles with deep origins in Asian or Black culture, e.g., “Can a White person wear a kimono?,” “Can a White person wear a durag?,” “Can White girls get braids?,” and “How does a White boy get waves?” (Figure 5).  [^5] https://www.youtube.com/watch?v=gSZ8U64EL-Q Additionally, we found instances of blackfishing, i.e., the practices of modifying one’s appearances physically or digitally or using terms and language patterns linked to Ebonics for racial commodification (Stevens, 2021, p.1). For example, we found autocompletes for White groups that included references to the makeup style of White people attempting to look like Blacks (“The White woman who turned Black”).
Figure 5: A YouTube video suggested by the autocomplete “Can a White person wear a durag?”.
In contrast, for Asians and Blacks, autocompletes related to cultural attire rarely reflected acts of cultural appropriation toward other cultures. Instead, the autocompletes for Blacks were related to changing their natural hairstyle (“How can a Black person get curly hair?,” “Can a Black man straighten his hair?”) or achieving a popular hairstyle or makeup (“How to do a Black person’s hair/makeup?,” “How to do a Black man bun?”).
The autocompletes for Asians often included the term “Asian Baby Girl (ABG)” (“What is an Asian baby girl?,” “What is an abg Asian baby?”). Asian Baby Girl is a slang term referring to “a young Asian woman who displays stereotypical traits, such as enjoying clubbing, wearing excessive makeup and tattoos, drinking bubble tea, wearing revealing clothes, etc.” (Wiktionary, 2023) Though the popularity of such language can subvert stereotypical images of Asian women as submissive or quiet, it also associates them with negative attributes, like aggression and violence (Wu, 2023). Thus, autocompletes about ABG could increase awareness about the diversity of Asian women but, in doing so, risk generating new negative stereotypes about them.
@@
Figure 6: A YouTube video suggested by the autocomplete “When an Asian is bad at math.”
@@
Ethnic Humor. Schutz defined ethnic humor as “humor directed at racial and nationality groups, denigrating alleged attributes of those groups” (Schutz, 1989, p.167). Ethnic stereotypes or shared beliefs toward specific ethnic groups are the basis of such humor (Boxman-Shabtai and Shifman, 2015). In our analysis, Asians were the target of the largest number of autocompletes in this category. Most of these autocompletes led to content created by the Chinese-Irish comedian Steven He O’Byrne (12.3 million subscribers). [^6] https://www.youtube.com/@StevenHe O’Byrne’s videos usually parody Asian dads and mock East-Asian parenting stereotypes. For instance, the video content corresponding to the autocomplete “When an Asian is bad at math” shows O’Byrne portraying Asian parents who have a high expectation for their children’s academic achievements and are harsh toward them when grades do not meet their standards (Figure 6). [^7] https://www.youtube.com/watch?v=6lMpjmPiTWE Other autocompletes, like “When an Asian is kidnapped,” “When an Asian is president,” and “When an Asian is your substitute,” lead to videos with similar stereotypical images of Asians. While these videos presumably aim to generate clicks and their associated revenue through humor, they reproduce stereotypes by portraying Asians as alien or different and endorsing the ‘model minority’ myth, which characterizes Asian Americans as being financially and educationally more advanced than other minority groups (Lee et al., 2008).
@@
Language. Language attributes (such as accents) tend to be understood as a strong indicator of ethnic categorization (Rakić et al., 2011). We found several autocompletes related to accents or speaking styles of each racial group. For example, the autocompletes for speaking styles included, “How to do a White girl voice,” “How to do a White person voice,” and “How to do a Black man voice.” Autocompletes for accents referenced the unique mode of pronunciation of certain racial groups (“Black man can’t pronounce,” “What does an Asian accent sound like?,” “All Latino accents,” “Woman with Spanish accent”). Videos resulting from such autocompletes often depicted individuals mimicking or mocking the voices and accents of other racial groups. This suggests that the multicultural aspects of U.S. language are still being leveraged to otherize specific groups.
@@
Relationships. Many autocomplete results in this category highlight anxieties about interracial relationships. For example, some autocompletes reflect curiosity or potential worries about relationships between White women and people with darker skin (“Do White girls like Brown guys?,” “Do White girls like Latinos?”). Similarly, many autocompletes referenced relationships between Black people and women from different cultures (“Can a Black man marry an Arab woman?”). These results often portray interracial relations as incongruous and reflect an unfavorable view of darker skin in the relationship contexts.
@@ 4.4 Social Equity
Diversity/Inclusion. Only Blacks and Hispanics had autocompletes associated with the issues of diversity and inclusion. Curiously, most of these autocompletes referred to the underrepresentation of racial minorities in popular culture (“Can a Black man be Captain America?,” “Is there a Hispanic Disney princess?”; Figure 7). [^8] https://www.youtube.com/watch?v=AQLUeshn3nE This reflects a cultural appetite for seeing racial diversity and encouraging inclusive casting decisions in mainstream media. While whitewashing has historically been salient in the entertainment industry, the percentage of White characters depicted on-screen has slightly declined in recent years, while the depiction of Asian characters has significantly increased (Smith et al., 2023). However, no notable changes have occurred in the prevalence of Black or Hispanic characters over the same period (Smith et al., 2023). Seemingly in response to these trends, YouTube autocompletes in this category reveal frustrations about persistent inequalities in the racial representation of Black and Hispanic character portrayals.
Figure 7: A YouTube video suggested by the autocomplete “Is there a Hispanic Disney princess?”.
Racial Justice. We observed autocompletes in this category only for the White and Black racial groups. First, we found autocompletes that referenced ‘White Lives Matter,’ a racist counter-movement that emerged in response to the ‘Black Lives Matter’ movement (“Do White Lives Matter one minute?,” “Do White Lives Matter shorts?”) (Figure 8). [^9] https://www.youtube.com/watch?v=GlMYUFDb8qg ‘White Lives Matter’ uses a variation of the Black Lives Matter core slogan to weaken the latter’s message about addressing unwarranted police violence against Blacks (Goodman et al., 2023). ‘White Lives Matter’ supporters have also been criticized by the Anti-Defamation League (Smith et al., 2023) for its associations with racism, and, in 2019, it was designated as a hate group by the Southern Poverty Law Center. Despite this controversy, the slogan ‘White Lives Matter’ appeared verbatim in YouTube autocompletes.
Despite video results for the query ‘White Lives Matter’ displaying diverse voices and criticisms on contested topics, text suggestions of YouTube autocomplete itself were solely generated based on the messages of ‘White Lives Matter,’ and no analogous autocompletes were associated with ‘Black Lives Matter’ messaging. This indicates that the autocompletes curated by YouTube do not always balance perspectives from opposing sides of an issue and, in this case, give short shrift to the side combatting racial disparities.
@@
Figure 8: A YouTube video suggested by the autocomplete “Do White Lives Matter one minute?”.
@@ 4.5 Manner
Aggression. Our results reference perceptions of Blacks and Asians as being aggressive. Blacks had the highest number of autocompletes in this category, which often described them as dangerous (“Pretend there is a Black man chasing you,” “Run like a Black man is chasing you”) or criminal (“A Black woman is accused of shoplifting”). This aligns with long-held negative stereotypes of Black violence and aggression (Gilens, 1996). We also found one autocomplete that suggests a desire to circumvent this stereotype: “Every Black man should have a latte.” While this does not relate to aggressive traits at face value, the top resulting video shows a Black stand-up comedian poking fun at the violence stereotype by recommending Black men hold a latte to appear ‘safe’ to law enforcement officials. Additionally, we found one aggression-related autocomplete for Asians: “Asian man goes crazy.” The top resulting video for this shows an Asian game streamer screaming violently while playing games. Other videos show Asian men engaging in gun violence. While such aggressive images are typically not associated with Asian men in the U.S., these videos depict Asian men as not always docile and potentially violent, especially in the context of online games.
Inappropriate. This category contains autocompletes that describe Black people as lazy or behaving inappropriately in social settings. One longstanding negative stereotype of Black people maintains that they are lazy, which is also posed as an explanation for other undesirable attributes, such as their disproportionate poverty (Gilens, 1996). The autocomplete “Black people can’t hear smoke detectors” depicts this stereotype; it led to videos that show Black people as being too lazy to change the low battery of smoke detectors even though their alarms kept sounding. Figure 9 presents a series of YouTube shorts that show how creators promote this stereotype.
Additionally, several autocompletes describe Black people as lacking the self-control needed to behave appropriately in social settings. For instance, this category included “Black man can’t hold his laugh” and “Black man can’t stop laughing.” The resulting videos show instances where a Black man cannot refrain from laughing in a serious situation where laughter is clearly inappropriate.
@@ 5 Discussion
We began this study with the goal of examining the information that YouTube provides to users seeking video content on race-related topics. We found that besides listing offerings that assuage users’ curiosities on these topics, YouTube plays a role in spreading racial stereotypes through its autocomplete function, a feature that incidentally cannot be turned off.
@@
Our analysis reveals that race-related autocompletes embody troubling stereotypes regarding Blacks’ personal hygiene, financial abilities, language use, aggression, and laziness; Whites’ athletic abilities; and Asians’ image as the high-achieving model minority. Notably, we observe that YouTube search has several unique characteristics that promote group stereotypes. For example, it amplifies distorted racial views that appear in mainstream entertainment media by reproducing titles or dialogues from existing media content (e.g., “White Men Can’t Jump,” “That is a White man’s luxury”). We also find that racial stereotypes on YouTube are often subtly delivered in the form of ethnic humor and memes. These genres and formats may make naive users embrace group stereotyping as a new and enjoyable form of entertainment rather than recognize its potential harms toward racial minorities, as evidenced by the millions of views and likes such content generates (e.g., O’Byrne’s videos).
@@
In addition to reinforcing racial stereotypes, our analysis of YouTube autocompletes also reflects tensions about interracial relationships, White appropriations of others’ cultural heritage, and Blacks’ and Hispanics’ concerns about their media representation. The norms of whiteness appear to prevail, with racial minorities struggling to combat structural inequalities (“Can a Black man be Captain America?”, “Is there a Hispanic Disney princess?”). These results serve as a lens through which to examine the current state of race in the U.S. Though some popular narratives maintain that we live in a ‘post-race’ society (Squires, 2014), the evidence of racial tensions we found in our analysis indicates otherwise. Despite a range of efforts to reduce racial discrimination, our study indicates the need for additional initiatives to promote racial harmony and resist the othering of racial minorities.
@@
Our results highlight ethical concerns about the potential ‘nudge effects’ of autocomplete suggestions on YouTube users. Graham (2023) noted that autocompletes can be biased in ways that might not be clear on an individual level but become apparent when evaluating a larger sample. Our analysis reveals evidence of similar aggregated discrimination in race-related queries on YouTube. For example, we find that while Black struggles to combat oppression are diminished, e.g., through the removal of references to “Black Lives Matter” in autocomplete suggestions, “White Lives Matter” continues to feature in these suggestions despite its known associations with hate groups.
@@
Prior studies have shown how biased autocomplete suggestions can impact users’ votes in political elections (Epstein et al., 2024), lead to hiring discrimination (Lin et al., 2024), and shift attitudes toward societal issues, e.g., support for fracking and legalizing the death penalty (Williams-Ceci et al., 2024). Thus, there is cause for concern that repeated exposure to racial stereotypes in YouTube autocompletes may adversely influence users’ attitudes and behaviors toward different racial groups. The black-box nature of such AI-based technologies is likely to prevent users from gaining access to information about how outputs were generated and how they could embed identity-based biases (Rai, 2020). Indeed, users may be totally unaware of AI biases and fail to notice the influence of these biases on their attitudes (Williams-Ceci et al., 2024). Future work can build upon this research by analyzing the cognitive processes that mediate how YouTube users are influenced by biases in autogenerated content on the site.
@@
Our findings indicate the need for moderation strategies for the ethical deployment of AI-based tools such as autocompletes. Prior research on moderation in search engine autocompletes has shown that platforms moderate derogatory stereotypes about many (although not all) key demographic categories (Leidinger and Rogers, 2023). Google even has an autocompletion moderation policy to prevent autocompletes “that might be violent, sexually-explicit, hateful, disparaging or dangerous” (Sullivan, 2020). Yet, as our findings show, YouTube, owned by the same parent company as Google, perpetuates racial stereotypes, highlighting the need for improved policy and corporate-wide enforcement. As Rogers (2023) points out, platforms like Google should take actions that exceed mere ‘quiet patches’ of infamous autocompletions highlighted by journalists and researchers on English language sites by ensuring that adequate remedies occur for all products (including Google Images and YouTube) and language offerings.
Beyond search autocompletes, AI mediation has been applied to a wide range of user-facing interfaces, such as auto-corrections and auto-responses. To counter discriminatory outputs in such features, we recommend that content curation infrastructures incorporate a regular review of automatically generated texts and suggested content, at least for the most frequently searched, culturally sensitive topics. We highlight that it is crucial to comprehensively consider the information-seeking contexts within which such race-based autocomplete queries are made and the user-uploaded content that may have shaped autocomplete results. Our analysis shows that in some cases (“Black women can’t buy a dress,” “White Lives Matter”), autocomplete suggestions themselves were problematic, but many resulting videos offered a more nuanced understanding and sometimes even countered pre-existing racial stereotypes. It appears that human moderators who understand subtle nuances, contexts, and the purpose of videos could be deployed to curate such algorithmic outputs. It would also be informative to analyze the relative influence of search autocomplete texts and their resulting videos on user attitudes and opinions.
Our research methods offer a blueprint for critically examining the identity-based biases in AI-generated outputs and shaping design and literacy measures to address those biases. More broadly, we need to adopt a transformative approach to systematically counter discrimination, both online and offline. To this end, content moderation and AI audit efforts must be paired with educational, cultural, and policy initiatives to combat such biases.

### v2

@@ 4.3 Culture
First, we found autocompletes that indicated appropriation of cultural practices, highlighting asymmetric power relations between racial groups. For example, several autocompletes related to White groups coopted indigenous Black cultures while belittling its native contexts (e.g., “white girl can dance african”). Other autocompletes revealed how White people could appropriate specific aspects of popular Asian culture, e.g., “can a white guy won [sic] an asian beauty pageant,” “can a white person become a kpop idol.” The keyword ‘Black’ elicited one autocomplete related to the adoption of Asian cultural elements (“can a black person become a kpop idol”).
Second, autocompletes associated with culturally specific attire and beauty standards also indicated instances of cultural appropriation, where members of the majority group coopted traditional garments, hairstyles, or makeup practices of other minority groups. For example, autocompletes related to garments or hairstyles for White groups included traditional styles with deep origins in Asian or Black culture, e.g., “can a white person wear a kimono,” “can a white person wear a durag,” “can white girls get braids,” and “can a white boy get waves” (Figure 5). [^5] https://www.youtube.com/watch?v=gSZ8U64EL-Q Additionally, we found instances of blackfishing, i.e., the practices of modifying one’s appearances physically or digitally or using terms and language patterns linked to Ebonics for racial commodification (Stevens, 2021, p.1). For example, we found autocompletes for White groups that included references to the makeup style of White people attempting to look like Blacks (“the white woman who turned black”).
Figure 5: A YouTube video suggested by the autocomplete “can a white person wear a durag.”
In contrast, for Asians and Blacks, autocompletes related to cultural attire rarely reflected acts of cultural appropriation toward other cultures. Instead, the autocompletes for Blacks were related to changing their natural hairstyle (“how can a black person get curly hair,” “can a black man straighten his hair”) or achieving a popular hairstyle or makeup (“how to do a black person’s hair,” “how to do a black person’s makeup,” “how to do a black man bun”).
The autocompletes for Asians often included the term “Asian Baby Girl (ABG)” (“what is an asian baby girl,” “what is an abg asian baby”). Asian Baby Girl is a slang term referring to “a young Asian woman who displays stereotypical traits, such as enjoying clubbing, wearing excessive makeup and tattoos, drinking bubble tea, wearing revealing clothes, etc” (Wiktionary, 2023). Though the popularity of such language can subvert stereotypical images of Asian women as submissive or quiet, it also associates them with negative attributes, like aggression and violence (Wu, 2023). Thus, autocompletes about ABG could increase awareness about the diversity of Asian women but, in doing so, risk generating new negative stereotypes about them.
@@
Figure 6: A YouTube video suggested by the autocomplete “when an asian is bad at maths.”
@@
Ethnic Humor. Schutz defined ethnic humor as “humor directed at racial and nationality groups, denigrating alleged attributes of those groups” (Schutz, 1989, p.167). Ethnic stereotypes or shared beliefs toward specific ethnic groups are the basis of such humor (Boxman-Shabtai and Shifman, 2015). In our analysis, Asians have the largest number of autocompletes in this category. Most of these autocompletes led to content created by the Chinese-Irish comedian Steven He O’Byrne (12.3 million subscribers). [^6] https://www.youtube.com/@StevenHe O’Byrne’s videos usually parody Asian dads and mock East-Asian parenting stereotypes. For instance, the video content corresponding to the autocomplete “when an asian is bad at maths” shows O’Byrne portraying Asian parents who have a high expectation for their children’s academic achievements and are harsh toward them when grades do not meet their standards (Figure 6). [^7] https://www.youtube.com/watch?v=6lMpjmPiTWE Other autocompletes, like “when an asian is kidnapped,” “when an asian is president,” and “when an asian is your substitute,” lead to videos with similar stereotypical images of Asians. While these videos presumably aim to generate clicks and their associated revenue through humor, they reproduce stereotypes by portraying Asians as alien or different and endorsing the ‘model minority’ myth, which characterizes Asian Americans as being financially and educationally more advanced than other minority groups (Lee et al., 2008).
@@
Language. Language attributes (such as accents) tend to be understood as a strong indicator of ethnic categorization (Rakić et al., 2011). We found several autocompletes related to accents or speaking styles of each racial group. For example, the autocompletes for speaking styles included, “how to do a white girl voice,” “how to do a white person voice,” and “how to do a black man voice.” Autocompletes for accents referenced the unique mode of pronunciation of certain racial groups (“black man can’t pronounce,” “what does an asian accent sound like,” “all latino accents,” “woman with spanish accent”). Videos resulting from such autocompletes often depicted individuals mimicking or mocking the voices and accents of other racial groups. This suggests that the multicultural aspects of U.S. languages are still being leveraged to otherize specific groups.
@@
Relationships. Many autocompletes in this category highlight anxieties about interracial relationships. For example, some autocompletes reflect curiosity or potential worries about relationships between White women and people with darker skin (“do white girls like brown guys,” “do white girls like latinos”). Similarly, many autocompletes referenced relationships between Black people and women from different cultures (“can a black man marry an arab woman”). These results often portray interracial relations as incongruous and reflect an unfavorable view of darker skin in the relationship contexts.
@@ 4.4 Social Equity
Diversity/Inclusion. Whites, Blacks, and Hispanics had autocompletes associated with the issues of diversity and inclusion. Curiously, most of these autocompletes referred to the underrepresentation of racial minorities in popular culture (“can a black man be captain america,” “is there a hispanic disney princess” Figure 7). [^8] https://www.youtube.com/watch?v=AQLUeshn3nE This reflects a cultural appetite for seeing racial diversity and encouraging inclusive casting decisions in mainstream media. While whitewashing has historically been salient in the entertainment industry, the percentage of White characters depicted on-screen has slightly declined in recent years, while the depiction of Asian characters has significantly increased (Smith et al., 2023). However, no notable changes have occurred in the prevalence of Black or Hispanic characters over the same period (Smith et al., 2023). Seemingly in response to these trends, YouTube autocompletes in this category reveal frustrations about persistent inequalities in the racial representation of Black and Hispanic character portrayals.
Figure 7: A YouTube video suggested by the autocomplete “is there a hispanic disney princess.”
Racial Justice. We observed autocompletes in this category only for the White and Black racial groups. First, we found autocompletes that referenced ‘White Lives Matter,’ a racist counter-movement that emerged in response to the ‘Black Lives Matter’ movement (“do white lives matter one minute,” “do white lives matter shorts”) (Figure 8). [^9] https://www.youtube.com/watch?v=GlMYUFDb8qg ‘White Lives Matter’ uses a variation of the Black Lives Matter core slogan to weaken the latter’s message about addressing unwarranted police violence against Blacks (Goodman et al., 2023). The slogan ‘White Lives Matter’ appeared verbatim in YouTube autocompletes. Despite video results for the query ‘White Lives Matter’ displaying diverse voices and criticisms on contested topics, text suggestions of YouTube autocomplete itself were solely generated based on the messages of ‘White Lives Matter,’ and no analogous autocompletes were associated with ‘Black Lives Matter’ messaging. This indicates that the autocompletes curated by YouTube do not always balance perspectives from opposing sides of an issue and, in this case, give short shrift to the side combatting racial disparities.
@@
Figure 8: A YouTube video suggested by the autocomplete “do white lives matter one minute.”
@@ 4.5 Manner
Aggression. Our results reference perceptions of Blacks and Asians as being aggressive. Blacks had the highest number of autocompletes, which often described them as dangerous (“pretend there is a black man chasing you,” “run like a black man is chasing you”) or criminal (“a black woman is accused of shoplifting”). This aligns with long-held negative stereotypes of Black violence and aggression (Gilens, 1996). We also found one autocomplete that suggests a desire to circumvent this stereotype: “every black man should have a latte.” While this does not relate to aggressive traits at face value, the top resulting video shows a Black stand-up comedian poking fun at the violence stereotype by recommending Black men hold a latte to appear ‘safe’ to law enforcement officials. Additionally, we found aggression-related autocompletes for Asians: “asian man goes crazy.” The top resulting video for this shows an Asian male game streamer screaming violently while playing games. Other videos show Asian men engaging in gun violence. While such aggressive images are typically not associated with Asian men in the U.S., these videos depict Asian men as not always docile and potentially violent.
Inappropriate Behavior. This category contains autocompletes that describe Black people as lazy or behaving inappropriately in social settings. One longstanding negative stereotype of Black people maintains that they are lazy, which is also posed as an explanation for other undesirable attributes, such as their disproportionate poverty (Gilens, 1996). The autocomplete “black people can’t hear smoke detectors” depicts this stereotype; it led to videos that show Black people as being too lazy to change the low battery of smoke detectors even though their alarms kept sounding. Figure 9 presents a series of YouTube shorts that show how creators promote this stereotype.
Several autocompletes describe Black people as lacking the self-control needed to behave appropriately in social settings. For instance, this category included “black man cant [sic] hold his laugh,” and “black man can’t stop laughing.” The resulting videos show instances where a Black man cannot refrain from laughing in a serious situation where laughter is clearly inappropriate.
@@ 5 Discussion
We began this study with the goal of examining the information that YouTube’s search feature provides to users seeking video content on race-related topics. We found that besides listing offerings that assuage users’ curiosities on these topics, YouTube plays a role in spreading racial stereotypes through its autocomplete function, a feature that incidentally cannot be turned off.
@@
Our analysis reveals that race-related autocompletes embody troubling stereotypes regarding Blacks’ personal hygiene, financial abilities, language use, aggression, and laziness; Whites’ athletic abilities; and Asians’ image as the high-achieving model minority. Notably, we observed that YouTube search amplifies distorted racial views that appear in mainstream entertainment media by reproducing titles or dialogues from existing media content (e.g., “white men can’t jump”). We posit that this can be attributed to YouTube’s participatory culture (Khan and Malik, 2022), where users actively contribute to content creation aimed at capturing popular attention and resulting in monetary gains. Thus, an interplay of mainstream media and amateur online cultures could be perpetuating racial stereotypes on the platform. Creators’ communal and commercial aspirations (Boxman-Shabtai, 2019) could also account for our finding that racial stereotypes on YouTube are often subtly delivered in the form of ethnic humor and memes. These genres and formats may make naive viewers embrace group stereotyping as a new and enjoyable form of entertainment rather than recognize its potential harm toward racial minorities, as evidenced by the millions of views and likes such content generates.
@@
In addition to reinforcing racial stereotypes, our analysis of YouTube autocompletes reflects tensions about interracial relationships, and Blacks’ and Hispanics’ concerns about their media representation. The norms of whiteness appear to prevail, with racial minorities struggling to combat structural inequalities (“can a black man be captain america,” “is there a hispanic disney princess”). These results serve as a lens through which to examine the current state of race in the U.S. Though some popular narratives maintain that we live in a ‘post-race’ society (Squires, 2014), the evidence of racial tensions we found in our analysis indicates otherwise. Despite a range of efforts to reduce racial discrimination, our study indicates the need for additional initiatives to promote racial harmony and resist the othering of racial minorities.
@@
Our results highlight ethical concerns about the potential ‘nudge effects’ of autocomplete suggestions on YouTube users. Graham (2023) noted that autocompletes can be biased in ways that might not be clear on an individual level but become apparent when evaluating a larger sample. Our analysis reveals evidence of similar aggregated discrimination in race-related queries on YouTube. For example, we find that while Black struggles to combat oppression are diminished, e.g., through the removal of references to “Black Lives Matter” in autocomplete suggestions, “White Lives Matter” continues to feature in these suggestions despite its known associations with hate groups. We also found traces of power imbalance through various ways of cultural appropriation reflected in YouTube autocompletes, showing how asymmetric racial values are embedded in YouTube’s technological configurations. Our unique methodological approach has helped us surface these biases, and it can also inform future analyses of autocomplete outputs, especially on social media and video-sharing platforms.
@@
Prior studies have shown how biased autocomplete suggestions can impact users’ votes in political elections (Epstein et al., 2024), lead to hiring discrimination (Lin et al., 2024), and shift attitudes toward societal issues, e.g., support for fracking and legalizing the death penalty (Williams-Ceci et al., 2024). Thus, there is cause for concern that repeated exposure to racial stereotypes in YouTube autocompletes may adversely influence users’ attitudes and behaviors toward different racial groups. The black-box nature of such AI-based technologies is likely to prevent users from gaining access to information about how outputs were generated and how they could embed identity-based biases (Rai, 2020). In fact, users may be totally unaware of AI biases and fail to notice the influence of these biases on their attitudes (Williams-Ceci et al., 2024). Future work can build upon this research by analyzing the cognitive processes that mediate how YouTube users are influenced by biases in autogenerated content on the site.
Indeed, previous research on Google Search shows that its autocomplete outputs are sourced not just from web content and users’ search activity but also from external digitization projects, especially Google Books (Graham, 2023). Lack of information about how autocompletes are produced allows platforms to avoid taking the full ethical responsibility for problematic suggestions, even in cases when they result from machine learning biases rather than prior search queries. Therefore, we call for greater transparency regarding autocomplete sources and algorithms to facilitate a more precise attribution of problematic suggestions.
@@
Our findings indicate the need for moderation strategies for the ethical deployment of AI-based tools such as autocompletes. Prior research on moderation in search engine autocompletes has shown that platforms moderate derogatory stereotypes about various demographic categories (Leidinger and Rogers, 2023). Google even has an autocompletion moderation policy to prevent autocompletes “that might be violent, sexually-explicit, hateful, disparaging or dangerous” (Sullivan, 2020). Yet, as our findings show, YouTube, owned by the same parent company as Google, perpetuates racial stereotypes, highlighting the need for improved policy and corporate-wide enforcement. As Rogers (2023) points out, platforms like Google should take actions that exceed mere ‘quiet patches’ of infamous autocompletions highlighted by journalists and researchers on English language sites by ensuring that adequate remedies occur for all their products and language offerings.
This study has some limitations. We observed an under-representation of Hispanics in our analysis of autocompletes. However, it is possible that this result is influenced by other terms referring to the “Hispanic” race (e.g., “Latino,” “Mexican”) being used colloquially more often than similar counterparts (e.g., “Caucasian,” “African-American”) for other races. Future research on autocomplete audits should also collect queries using such colloquial terms.
A Black person can
can’t date a black woman until my grandpa died Does a Black woman
A Black woman can
can’t keep a black man down Does a Black man
Can a Black man
how often should a black man wash his hair Should a Black
Should a Black man
A Black man should
how often should a black woman wash her hair Should a Black
Should a Black woman
A Black woman should
how often should a black man wash his beard Should a Black man
black woman can’t buy a dress A Black woman can
a black man can’t have a suitcase A Black man can
run like a black man is chasing you A Black man is
every black man should have a latte A Black man should
black man can’t pronounce Black men can
black man can’t dance Black men can
black man can’t play basketball Black men can
black man can’t say beginning Black men can
black man cannot own g wagon Black men can
black man can’t stop laughing Black men can
black man cant [sic] hold his laugh Black men can
black people can’t marry meme Black people can
black people can’t marry white people Black people can
black people can’t name one african country Black people can
black people can’t hear smoke detectors Black people can
Asian what does an asian accent sound like Does an Asian
what is an asian baby girl Is an Asian
what is an abg asian baby Is an Asian girl
when an asian is bad at maths An Asian is
when an asian is kidnapped An Asian is
when an asian is president An Asian is
when an asian is your substitute An Asian is
asian man goes crazy An Asian man is
Hispanic is there a hispanic disney princess Is a Hispanic
all latino accents A Hispanic should
woman with spanish accent A Hispanic girl should
Table 4: This table lists autocomplete suggestions and the input queries they derived from for cases where the suggestion did not contain the input query. This highlights that YouTube’s autocomplete generation does not strictly follow the exact syntactic patterns of input queries.